OpenStudy (anonymous):

reduce y=x^2+4x+3 to standard form y-b+A(x-a)^2 or x-a=A(y-b)^2

OpenStudy (anonymous):

oh yeah and im somehow supposed to complete the square too..

OpenStudy (dumbcow):

use completing the square subtract 3 on both sides y-3 = x^2 +4x Take the x coefficient, halve it ,then square it, then add to both sides 4/2 = 2 -> 2^2 = 4 y-3 +4 = x^2 +4x +4 factor, notice right side is now a perfect square y+1 = (x+2)^2

OpenStudy (dumbcow):

This gives an upward opening parabola with vertex at (-2,-1)

OpenStudy (anonymous):

awsum thanks!